### v1

# Accessibility of Sparse Sets

Abstract.
A set $D\subseteq\mathbb{N}$ is called $r$-large if every $r$-coloring of $\mathbb{N}$ admits arbitrarily long monochromatic arithmetic progressions $a,a+d,...,a+(k-1)d$ with gap $d\in D$. Closely related to largeness is accessibility; a set $D\subseteq\mathbb{N}$ is called $r$-accessible if every $r$-coloring of $\mathbb{N}$ admits arbitrarily long monochromatic sequences $x_{1},x_{2},...,x_{k}$ with $x_{i+1}-x_{i}\in D$. It is known that if $D\subseteq\mathbb{N}$ is $2$-large, then the gaps between elements in $D$ cannot grow exponentially. In this paper, we show that if $D$ is $2$-accessible, then the gaps between elements in $D$ cannot grow much faster than exponentially.
1. Introduction
An $r$-coloring of a set $A$ is a function $\chi:A\to[r]$. We say a subset $A^{\prime}\subseteq A$ is monochromatic under $\chi$ if $\chi$ is constant over $A^{\prime}$. One of the main goals in the study of Ramsey Theory is to find “order” in seemingly “random” structures. For example, Van der Waerden’s Theorem tells us that given any $r$-coloring of the positive integers, there will exist arbitrarily long monochromatic arithmetic progressions. The theorem places no requirement on the gap (common difference), $d$, of the arithmetic progression – it can be any natural number. With this in mind, in [1], the authors introduce the notion of largeness.
Definition 1.1.
A set $D\subseteq\mathbb{N}$ is called $r$-large if, for every $r$-coloring of $\mathbb{N}$, there exist arbitrarily long monochromatic arithmetic progressions with gap $d\in D$. If $D$ is $r$-large for every $r\in\mathbb{N}$, we say $D$ is large.
In an attempt to learn more about largeness, in [8], the authors study some Ramsey properties of a larger family of sequences including arithmetic progressions.
Definition 1.2.
Let $D\subseteq\mathbb{N}$. A $k$-term sequence of positive integers $x_{1},x_{2},...,x_{k}$ where $x_{i+1}-x_{i}\in D$ is called a $k$-term $D$-diffsequence.
Definition 1.3.
A set $D\subseteq\mathbb{N}$ is called $r$-accessible if, for every $r$-coloring of $\mathbb{N}$, there exist arbitrarily long monochromatic sequences $x_{1},x_{2},...,x_{k}$ with $x_{i+1}-x_{i}\in D$, that is, there exist monochromatic $k$-term $D$-diffsequences for every $k\geq 1$. Similarly, if $D$ is $r$-accessible for every $r\geq 1$, we say $D$ is accessible.
Equivalently, by a compactness argument, a set $D\subseteq\mathbb{N}$ is $r$-accessible if, for every $k\geq 1$, there exists a least positive integer $\Delta=\Delta(D,k;r)$ such that whenever $[1,\Delta]$ is $r$-colored, there exists a monochromatic $k$-term $D$-diffsequence. The same is true for $r$-largeness. Given a set $D\subseteq\mathbb{N}$, we define the degree of accessibility of $D$, $\text{doa}(D)$, to be the greatest positive integer $r$ such that $D$ is $r$-accessible. If $D$ is $r$-accessible for every $r\in\mathbb{N}$, we write $\text{doa}(D)=\infty.$ Finally, note that it follows immediately from the definitions of largeness and accessibility that any set $D\subseteq\mathbb{N}$ which is $r$-large is automatically $r$-accessible.
It is a consequence of the Polynomial Van der Waerden Theorem [10] that given any polynomial $p\in\mathbb{Q}[x]$ with positive leading term and zero constant term, the set $\{p(n):n\in\mathbb{N}\}\cap\mathbb{N}$ is large (and thus accessible). This gives us a large class of sets which are large. For example, this result tells us that the set $\{n^{k}:n\in\mathbb{N}\}$ is large for every $k\in\mathbb{N}$. Note that the preceding set has natural density 00 and is large, while the set $\{2k+1:k\geq 0\}$ of odd positive integers has natural density $1/2$ and is not even $2$-accessible (color the even positive integers blue and the odd positive integers red). This seems to suggest that the natural density of a set does not give us much information in regards to largeness and accessibility. However, in [1] the authors show that if $D=\{d_{1},d_{2},d_{3},...,\}$ is a set of positive integers satisfying $d_{i+1}\geq 3d_{i}$ for all $i\geq 1$, then $D$ is not $2$-large. In ([2], Question 5.4), the author asks if a similar result holds for $2$-accessibility. Namely, Does there exist an absolute constant $C\geq 2$ such that there is no $2$-accessible set $D=\{d_{1},d_{2},d_{3},...\}$ satisfying $d_{i+1}>Cd_{i}$ for all $i\in\mathbb{N}$? If yes, what is the smallest such $C$? We prove the following theorem:
Theorem 1.4.
Let $r\geq 2$ and let $D=\{d_{1},d_{2},d_{3},...\}$ be a set of positive integers. If there exists some $\delta>0$ such that
$d_{i+1}\geq\left(\frac{2r-1}{r-1}+\delta\right)d_{i}$
for all $i\geq 1$, then $\text{doa}(D)\leq r-1$.
Letting $r=2$, we immediately obtain the following corollary.
Corollary 1.5.
Let $D=\{d_{1},d_{2},d_{3},...\}$ be a set of positive integers. If there exists some $\delta>0$ such that
$d_{i+1}\geq\left(3+\delta\right)d_{i}$
for all $i\geq 1$, then $D$ is not $2$-accessible.
Corollary 1.5 is strikingly similar to the above mentioned growth rate condition for $2$-largeness; the only difference being the added $\delta>0$. The $3$ found in both results seems to be a coincidence, however. As we will see in section 4, the condition $d_{i+1}\geq 3d_{i}$ in the $2$-largeness result can be lowered to $d_{i+1}\geq(1+\delta)d_{i}$ where $\delta>0$ is any positive constant [4]. On the other hand, the set $\{2^{n}:n\geq 0\}$ is $2$-accessible and has $d_{i+1}\geq 2d_{i}$.
In [6], the authors ask about the degree of accessibility of the set of Fibonacci numbers $F=\{1,2,3,5,8,13,...\}$ and show that $2\leq\text{doa}(F)\leq 5$. Recently, in [11], the author lowers the upper bound, showing $\text{doa}(F)\leq 3$. The proof uses various properties of the Fibonacci sequence to construct a $4$-coloring of $\mathbb{N}$ that avoids arbitrarily long monochromatic $F$-diffsequences. In this paper, we arrive at the same upper bound of $\text{doa}(F)\leq 3$ using only the growth rate of the Fibonacci sequence.
2. Some Background and Basic Results
First, we note that if $D$ is a finite set, then $D$ is neither $2$-large nor $2$-accessible. To see this, let $n=\max(D)$. Then, the $2$-coloring $\chi:\mathbb{N}\to\{1,2\}$ such that
$\chi(x)=\begin{cases}1&\text{if }1\leq(x\bmod{2n})\leq n\\ 2&\text{otherwise }\end{cases}$
admits no monochromatic $(n+1)$-term $D$-$AP$ and no $(n+1)$-term $D$-diffsequence.
The following result gives us a quite powerful necessary condition for a set $D$ to be $2$-large. A proof can be found in [1].
Theorem 2.1 ([1], Theorem 2.1).
If $D\in\mathbb{N}$ is $2$-large then $D$ contains a multiple of $m$ for every $m\in\mathbb{N}$.
We also have the previously mentioned growth rate condition for $2$-largeness.
Theorem 2.2 ([1], Theorem 2.2).
Let $A=\{a_{n}\}_{n=1}^{\infty}$ be a sequence of positive integers where $a_{n}\geq 3a_{n-1}$ for all $n\geq 2$. Then $A$ is not $2$-large.
The following tells us that if $D=D_{1}\cup D_{2}$ is large, then at least one of $D_{1}$ or $D_{2}$ is also large. A inductive argument naturally generalizes to $D=\bigcup_{i=1}^{n}D_{i}$ (let $D_{1}=\bigcup_{i=1}^{n-1}D_{i}$ and $D_{2}=D_{n}$).
Theorem 2.3 ([1], Theorem 2.4).
If $D=D_{1}\cup D_{2}$ is large, then either $D_{1}$ or $D_{2}$ is large.
Probably the most general characterization of large sets is a consequence of the Polynomial Van der Waerden Theorem:
Theorem 2.4 ([1], Theorem 3.1).
let $p$ be a polynomial with integer coefficients, a positive leading term, and $p(0)=0$. Then, $\mathbb{N}\cap\{p(n):n\geq 1\}$ is large.
Remark: In the above theorem, it states that the polynomial need have integer coefficients. This is not necessary, the polynomial need only have rational coefficients. To see why, let $p(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+...+a_{1}x$ be a polynomial with $a_{n},a_{n-1},...,a_{1}\in\mathbb{Q}$. Then, we can write
$a_{n}=\frac{p_{n}}{q_{n}},a_{n-1}=\frac{p_{n-1}}{q_{n-1}},...,a_{1}=\frac{p_{1}}{q_{1}}$
where $\gcd(p_{i},q_{i})=1$. Let $l=\text{lcm}(q_{1},q_{2},...,q_{n}).$ Then, $q(x)=p(lx)$ is a polynomial with integer coefficients and we can apply Theorem 2.4 to $q(x)$.
There are far fewer known results concerning only accessibility. The following is shown in [5]:
Theorem 2.5 ([5], Lemma 2.1).
Let $c\geq 0$ and $r\geq 2$, and let $D\subseteq\mathbb{N}$. If every $(r-1)$-coloring of $D$ yields arbitrarily long monochromatic $(D+c)$-diffsequences, then $D+c$ is $r$-accessible.
Most useful is the case when $r=2$ and $c=0$, in which case we have:
Corollary 2.6.
Let $D\subseteq\mathbb{N}$. If $D$ itself contains arbitrarily long $D$-diffsequences, then $D$ is $2$-accessible.
This corollary gives us the lower bound of $2$ for the degree of accessibility of the Fibonacci numbers $F$. Since $F=\{f_{n}:n\geq 1\}$ and $f_{n}=f_{n-1}+f_{n-2}$, we have that $f_{n}-f_{n-1}=f_{n-2}$ for all $n\geq 3$. Thus, the set of Fibonacci numbers contain infinitely long Fibonacci diffsequences, and are thus $2$-accessible by Corollary 2.6.
Given some subset $D\subseteq\mathbb{N}$, it is often useful to work with a subset of $D$, as opposed to $D$ itself. This next theorem is a generalization of a result that appears in [11] and [7].
Theorem 2.7.
Let $A\subseteq\mathbb{N}$ and let $p\in\mathbb{N}$. Define $B:=\left\{a_{i}/p:a_{i}\in A\right\}\cap\mathbb{N}$ and define $C:=\{a_{i}\in A:a_{i}\equiv 0\bmod{p}\}$. Then, the following holds:
(i) If $A$ is $rp$-accessible, then $B$ is $r$-accessible.
(ii) If $A$ is $rp$-accessible, then $C$ is $r$-accessible.
Proof.
In both cases, we will prove the contrapositive. First, assume that $B$ is not $r$-accessible. Then, there exists $k\in\mathbb{N}$ and a coloring $\chi:\mathbb{N}\to[r]$ such that $\chi$ admits no monochromatic $k$-term $B$-diffsequence. Now, assume for the sake of contradiction that $A$ is $rp$-accessible. Define the following $rp$-coloring $\chi_{1}:\mathbb{N}\to[p]\times[r]$ such that
$\chi_{1}(x)=\left(x\bmod{p},\chi\left(\frac{x-x\bmod{p}}{p}+1\right)\right).$
Since $A$ is $rp$-accessible, there exists a monochromatic $k$-term $A$-diffsequence $X=\{x_{1},x_{2},...,x_{k}\}$ under $\chi_{1}$. By definition of $\chi_{1}$, every element in $X$ is congruent to $c\bmod{p}$ for some $c\in\{0,1,...,p-1\}$. But then,
$\chi\left(\frac{x_{1}-c}{p}+1\right)=\chi\left(\frac{x_{2}-c}{p}+1\right)=\cdots=\chi\left(\frac{x_{k}-c}{p}+1\right)$
and
$\left(\frac{x_{j+1}-c}{p}+1\right)-\left(\frac{x_{j}-c}{p}+1\right)=\left(\frac{x_{j+1}-x_{j}}{p}\right)\in B$
for all $j\in\{1,2,...,k-1\}$, a contradiction.
Now, assume that $C$ is not $r$-accessible. Then, there exists $k\in\mathbb{N}$ and a coloring $\chi:\mathbb{N}\to[r]$ such that $\chi$ admits no monochromatic $k$-term $C$-diffsequence. Again, assume for the sake of contradiction that $A$ is $rp$-accessible. Define the following $rp$-coloring $\chi_{1}:\mathbb{N}\to[p]\times[r]$ such that
$\chi_{1}(x)=\left(x\bmod{p},\chi(x)\right).$
Since $A$ is $rp$-accessible, there exists a monochromatic $k$-term $A$-diffsequence $X=\{x_{1},x_{2},...,x_{k}\}$ under $\chi_{1}$. By definition of $\chi_{1}$, every element in $X$ is congruent to $c\bmod{p}$ for some $c\in\{0,1,...,p-1\}$. But then,
$\chi(x_{1})=\chi(x_{2})=\cdots=\chi(x_{k})$
and $x_{j+1}-x_{j}\in A$ and $x_{j+1}-x_{j}\equiv 0\bmod{p}$, so $x_{j+1}-x_{j}\in C$ for all $j\in\{1,2,...,k-1\}$, a contradiction.
Theorem 2.7 will be useful for showing $\text{doa}(F)\leq 3$. By letting $F_{E}=\{f_{3n}:n\geq 1\}=\{2,8,34,...\}$, we see that $F_{E}=\{f\in F:f\equiv 0\bmod{2}\}$. Thus, if we can show $\text{doa}(F_{E})=1$, we will have $\text{doa}(F)\leq 3$.
3. Proof of Theorem 1.4 and some Consequences
We begin by giving a proof of our main result, Theorem 1.4. We define $\{x\}:=x-\lfloor x\rfloor$ to be the fractional part of $x\in\mathbb{R}$. The first part of the proof uses the fact that $d_{i+1}\geq\left(\frac{2r-1}{r-1}+\delta\right)d_{i}$ for all $i\geq 1$ to construct an $\alpha\in\mathbb{R}$ satisfying $\{\alpha d_{i}\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $i\geq 1$, where $\varepsilon=\varepsilon(\delta)>0$ is a fixed positive constant. The second part shows that this condition on the fractional parts $\{\alpha d_{i}\}$ is sufficient to conclude $D$ is not $r$-accessible.
Proof of Theorem 1.4.
Let $D=\{d_{1},d_{2},d_{3},...\}$ be a set of positive integers such that there exists $\delta>0$ satisfying
$d_{i+1}\geq\left(\frac{2r-1}{r-1}+\delta\right)d_{i}$
for all $i\geq 1$. We begin by constructing a sequence of nested intervals $I_{1}\supseteq I_{2}\supseteq I_{3}\supseteq\cdots$ such that
$I_{n}=\left[\frac{z_{n}+\varepsilon}{d_{n}},\frac{rz_{n}+(r-1)}{rd_{n}}\right],$
where $z_{n}\in\mathbb{N}\cup\{0\}$ and $\varepsilon>0$ is fixed. Set
$\varepsilon=\frac{\delta(r-1)^{2}}{r\left(2r-1+\delta(r-1)\right)}$
and define $I_{1}=\left[\frac{\varepsilon}{d_{1}},\frac{r-1}{rd_{1}}\right].$ Assume we have defined $I_{1}\supseteq I_{2}\supseteq\cdots\supseteq I_{k}$ as above. Thus, we have
$I_{k}=\left[\frac{z_{k}+\varepsilon}{d_{k}},\frac{rz_{k}+(r-1)}{rd_{k}}\right].$
Note that $I_{k}=I^{1}_{k}\cup I^{2}_{k}$ where
$I^{1}_{k}=\left[\frac{z_{k}+\varepsilon}{d_{k}},\frac{z_{k}+\varepsilon}{d_{k}}+\frac{1}{d_{k+1}}\right]$
and
$I^{2}_{k}=\left[\frac{z_{k}+\varepsilon}{d_{k}}+\frac{1}{d_{k+1}},\frac{rz_{k}+(r-1)}{rd_{k}}\right].$
Since $|I^{1}_{k}|\geq\frac{1}{d_{k+1}}$, there must exist $z_{k+1}\in\mathbb{N}\cup\{0\}$ such that $\frac{z_{k+1}}{d_{k+1}}\in I^{1}_{k}$. We claim that
$I_{k+1}=\left[\frac{z_{k+1}+\varepsilon}{d_{k+1}},\frac{rz_{k+1}+(r-1)}{rd_{k+1}}\right]\subseteq I_{k}.$
Since $\frac{z_{k+1}}{d_{k+1}}\in I_{k}^{1}$ and $I_{k}=I_{k}^{1}\cup I_{k}^{2}$, the above inclusion will hold if $|I^{2}_{k}|\geq\frac{r-1}{rd_{k+1}}.$ By construction,
$|I^{2}_{k}|=\frac{r(1-\varepsilon)-1}{rd_{k}}-\frac{1}{d_{k+1}}.$
Now,
$\frac{r(1-\varepsilon)-1}{rd_{k}}-\frac{1}{d_{k+1}}\geq\frac{r-1}{rd_{k+1}}$ $\Longleftrightarrow\frac{r(1-\varepsilon)-1}{d_{k}}\geq\frac{2r-1}{d_{k+1}}$
$\Longleftrightarrow\left(r(1-\varepsilon)-1\right)d_{k+1}\geq\left(2r-1\right)d_{k}$
$\Longleftrightarrow d_{k+1}\geq\left(\frac{2r-1}{r(1-\varepsilon)-1}\right)d_{k}$
$\Longleftrightarrow d_{k+1}\geq\left(\frac{2r-1}{r-1}+\delta\right)d_{k}$
which is true by our choice of $\varepsilon.$ Now, since $|I_{n}|\to 0$ as $n\to\infty$, we have $\bigcap_{n=1}^{\infty}I_{n}=\{\alpha\}$. By construction, this $\alpha\in\mathbb{R}$ satisfies $\{\alpha d_{i}\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $i\geq 1$.
Consider the $r$-coloring $\chi:\mathbb{N}\to\{1,2,...,r\}$ defined by
$\chi(x)=\begin{cases}1&\text{if }\{\alpha x\}\in\left[0,\frac{1}{r}\right)\\ 2&\text{if }\{\alpha x\}\in\left[\frac{1}{r},\frac{2}{r}\right)\\ \vdots&\\ r&\text{if }\{\alpha x\}\in\left[\frac{r-1}{r},1\right)\end{cases}$
We claim this coloring avoids monochromatic $\left(\lceil 1/r\varepsilon\rceil+1\right)$-term $D$-diffsequences. To see this, assume, for the sake of contradiction, that $X=\{x_{1},x_{2},...,x_{k}\}$ where $k=\left(\lceil 1/r\varepsilon\rceil+1\right)$ is monochromatic, say color $c$. Then, $\{\alpha x_{1}\}\in\left[\frac{c-1}{r},\frac{c}{r}\right)$. We claim
$\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}$
for all $11\leq n\leq k-$ 1. To see why, assume, for the sake of contradiction, that $\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}-1$. By our assumption that $\chi(X)=c$, we have $\frac{c-1}{r}\leq\{\alpha x_{n}\}<\frac{c}{r}$ and $\varepsilon\leq\{\alpha(x_{n+1}-x_{n})\}\leq\frac{r-1}{r}$ for all $11\leq n\leq k-$ 1. Thus, we have
$\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}-1<\frac{r-1}{r}+\frac{c}{r}-1=\frac{c-1}{r},$
a contradiction with $\chi(x_{m})=c$. Iterating this fact, we obtain
$\{\alpha x_{k}\}=\{\alpha(x_{k}-x_{k-1})\}+\{\alpha(x_{k-1}-x_{k-2})\}+\cdots+\{\alpha(x_{2}-x_{1})\}+\{\alpha x_{1}\}\geq(k-1)\varepsilon+\frac{c-1}{r}\geq\frac{c}{r}.$
Thus, $\chi(x_{k})\not=c$, a contradiction. It follows that $\text{doa}(D)\leq r-1.$
Remark: From the proof of Theorem 1.4, we obtain a new way of showing a set $D$ is not $r$-accessible. Namely, if there exists some $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\{\alpha d\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $d\in D$, then $D$ is not $r$-accessible.
Using the fact that the set of even Fibonacci numbers satisfy the recurrence $e_{n}=4e_{n-1}+e_{n-2}$ for all $n\geq 3$ (with $e_{1}=2$ and $e_{2}=8$), it follows by Corollary 1.5 that $\text{doa}(F_{E})=1$. Thus, by the comment following the proof of Theorem 2.7, we have $\text{doa}(F)\leq 3$. However, as seen in the following corollary, we can actually calculate an explicit upper bound for the longest monochromatic $F$-diffsequence:
Corollary 3.1.
Let $F_{E}=\{f_{3n}:n\geq 1\}=\{2,8,34,...\}$ be the set of even Fibonacci numbers. Then $\text{doa}(F_{E})=1$.
Proof.
Let $\phi=(1+\sqrt{5})/2$, $\varphi=(1-\sqrt{5})/2$, and $\alpha=1/\sqrt{5}.$ Then,
$f_{n}=\alpha\phi^{n}-\alpha\varphi^{n}$
for all $n\geq 0$. We note that
$\phi f_{n}=f_{n+1}-\varphi^{n}.$
Thus,
$(1+\phi)f_{n}=f_{n}+f_{n+1}-\varphi^{n}.$
We will use the fact that $f_{3n}+f_{3n+1}\equiv 1\bmod{4}.$ This is true for $n\in\{0,1\}$. Now, assume that $k\geq 2$ and $f_{3l}+f_{3l+1}\equiv 1\bmod{4}$ for all $l<k$. Then, $f_{3k}+f_{3k+1}\equiv f_{3(k-2)}+f_{3(k-2)+1}\equiv 1\bmod{4}$ since $f_{n}\equiv f_{n+6}\bmod{4}$ (the Fibonacci sequence is periodic modulo $m$ for every $m$, when $m=4$, the period is $6$). Hence, it follows that
$\left\{\frac{f_{3n}+f_{3n+1}}{4}\right\}=\frac{1}{4}$
for $n\geq 1$. For $n\geq 1$, we also have
$-0.04<-\frac{1}{4}\cdot\left(\frac{1-\sqrt{5}}{2}\right)^{4}\leq-\frac{\varphi^{3n}}{4}\leq-\frac{1}{4}\cdot\left(\frac{1-\sqrt{5}}{2}\right)^{3}<0.06$
Thus,
$0.310.21<\left\{(1+\phi)\cdot\frac{f_{3n}}{4}\right\}<$ 0.31
for all $n\geq 1$ and it follows by the remark following the proof of Theorem 1.4 that $\text{doa}(F_{E})=1$. In fact, the longest monochromatic $F_{E}$-diffsequence under $\chi$ is bounded above by $\lceil 1/2(0.21)\rceil+1=4$.
4. A Small Digression Into Largeness
It was mentioned in the introduction that the $d_{i+1}\geq 3d_{i}$ condition in Theorem 2.2 could be weakened to $d_{i+1}\geq(1+\delta)d_{i}$, where $\delta$ is any positive constant [4]. In this section, we prove a slight generalization, namely, we show that if the gaps between elements in $D$ eventually grow exponentially, then $D$ is not $2$-large:
Theorem 4.1.
[…]
$d_{n+1}\geq(1+\delta)d_{i}$
for all $n\geq N,$ then $D$ is not $2$-large.
We define $||x||:=\min(\{x\},1-\{x\})$, that is, the distance from $x$ to the nearest integer. We require the following result on sequences with exponential growth.
Theorem 4.2 ([9]).
Let $\{d_{n}\}_{n=1}^{\infty}$ be a sequence of positive integers such that
$d_{i+1}\geq(1+\delta)d_{i}$
for all $i\geq 1$. Then, there exists $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\|\alpha d_{n}\|\geq\varepsilon$ for all $n\geq 1$.
The main component of the proof of Theorem 4.1 relies on the following result. A proof appears in ([4], Theorem 2). The crux of our argument is practically the same, however, we use slightly different notation in our proof.
Theorem 4.3.
Let $D\subseteq\mathbb{N}$. If there exists an $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\|\alpha d\|\geq\varepsilon$ for all $d\in D$, then $D$ is not $2$-large.
Proof.
Since $||\alpha d||\geq\varepsilon$ for all $d\in D$, we have $\{\alpha d\}\in[\varepsilon,1-\varepsilon]$ for all $d\in D$. Now, we define the $2$-coloring $\chi:\mathbb{N}\to\{1,2\}$ by
$\chi(x)=\begin{cases}1:\{\alpha x\}\in\left[0,\frac{1}{2}\right)\\ 2:\{\alpha x\}\in\left[\frac{1}{2},1\right).\end{cases}$
Since $\{\alpha d\}\in[\varepsilon,1-\varepsilon]$ for all $d$, for any monochromatic arithmetic progression with $d\in D$ under $\chi$, we would have either $\{\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$ or $\{\alpha d\}\in\left[\frac{1}{2},1-\varepsilon\right]$. Now, if $\{\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$, then the same argument used in the latter part of Theorem 1.4 shows us that we cannot have a monochromatic $(\lfloor 1/2\varepsilon\rfloor+1)$-term arithmetic progression with gap $d$. If $\{\alpha d\}\in\left[\frac{1}{2},1-\varepsilon\right]$, then $\{-\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$. Thus, the same argument used backwards shows us that we cannot have a monochromatic $(\lfloor 1/2\varepsilon\rfloor+1)$-term arithmetic progression with gap $d$. Thus, $D$ is not $2$-large. ∎
Remark: Theorem 4.3 implies Theorem 2.1 since if $D\subseteq\mathbb{N}$ contains no multiple of $m\in\mathbb{N}$, then $||\frac{1}{m}d||\geq\frac{1}{m}$ for all $d\in D$.
With Theorems 4.2 and 4.3, we are now equipped to prove Theorem 4.1:
Proof of Theorem 4.1.
Let $D=\{d_{1},d_{2},d_{3},...\}$ satisfy the conditions of Theorem 4.1. If $D$ contains no multiple of $m\in\mathbb{N}$ for some $m$, then $D$ is not $2$-large by Theorem 2.1. Now, if $D$ contains a multiple of $m$ for every $m\in\mathbb{N}$, then any $2$-coloring of $\mathbb{N}$ that is eventually periodic will admit infinitely long monochromatic $D$-$AP$’s. Now, if $D$ contains a multiple of $m\in\mathbb{N}$ for every $m$, then $D^{*}=\{d_{N+1},d_{N+2},d_{N+3},...\}$ does as well ($D$ must contain a multiple of $m^{\prime}=m\cdot\max\{d_{1},d_{2},...,d_{N}\}$, and this multiple must be in $D^{*}$). Thus, $D^{*}$ satisfies the conditions for Theorem 4.2, so there exists an $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\|\alpha d\|\geq\varepsilon$ for all $d\in D^{*}$. Since $D^{*}$ contains a multiple of $m$ for every $m\in\mathbb{N}$, $\alpha$ must be irrational. Thus, $\|\alpha n\|\not=0$ for every $n\in\mathbb{N}$. Take
$\varepsilon^{\prime}=\min_{d\in D\setminus D^{*}}\{\varepsilon,\|\alpha d\|\}.$
Then, $\|\alpha d\|\geq\varepsilon^{\prime}$ for all $d\in D$, and $D$ is not $2$-large by Theorem 4.3.
It follows immediately from Theorem 4.1 that the set of Fibonacci numbers $F$ are not $2$-large. However, using only Theorem 4.3 and some properties of the Fibonacci sequence we can obtain an upper bound on the longest monochromatic arithmetic progression with gap $f\in F$:
Corollary 4.4.
The set $F=\{1,2,3,5,8,...\}$ of Fibonacci numbers is not $2$-large.
Proof.
Let $\phi=(1+\sqrt{5})/2$, $\varphi=(1-\sqrt{5})/2$, and $\alpha=1/\sqrt{5}.$ Then,
$f_{n}=\alpha\phi^{n}-\alpha\varphi^{n}$
for all $n\geq 0$. We note that
$\phi^{-1}f_{n}=f_{n-1}-\varphi^{n}$
and
$\phi f_{n}=f_{n+1}-\varphi^{n}.$
Thus,
$(\phi^{-1}+\phi)f_{n}=f_{n-1}+f_{n+1}-2\varphi^{n}.$
We will use the fact that $f_{n-1}+f_{n+1}\not\equiv 0\bmod{8}$ for all $n\geq 1$. This can be easily verified by checking the first few cases and then noting that $f_{n}\equiv f_{n+12}\bmod{8}.$ Direct computation gives us
$\left\{(\phi^{-1}+\phi)\cdot\frac{f_{1}}{8}\right\}=\left\{(\phi^{-1}+\phi)\cdot\frac{f_{2}}{8}\right\}\approx 0.28$
Then, for $n\geq 3$, we note that
$-0.04<-\frac{1}{4}\cdot\left(\frac{1-\sqrt{5}}{2}\right)^{4}\leq-\frac{\varphi^{n}}{4}\leq-\frac{1}{4}\cdot\left(\frac{1-\sqrt{5}}{2}\right)^{3}<0.06.$
Since
$\left\{\frac{f_{n-1}+f_{n+1}}{8}\right\}\in\left\{\frac{1}{8},\frac{2}{8},...,\frac{7}{8}\right\},$
we have
$0.9350.085<\left\{\left((\phi^{-1}+\phi)\cdot\frac{f_{n}}{8}\right)\right\}<$ 0.935 .
Thus, $\left\|\left(\frac{\phi^{-1}+\phi}{8}\right)f\right\|\geq 0.06$ for all $f\in F$ and $F$ is not $2$-large by Theorem 4.3. In fact, the longest monochromatic arithmetic progression with gap $f\in F$ under $\chi$ is bounded above by $\lceil 1/2(0.06)\rceil+1=10$.
5. Some Concluding Remarks
The condition $d_{i+1}\geq(1+\delta)d_{i}$ in Theorem 4.1 is tight. On the other hand, we are not sure whether the condition $d_{i+1}\geq(3+\delta)d_{i}$ in Corollary 1.5 is the best possible. We hoped to weaken the condition to $d_{i+1}\geq(2+\delta)d_{i}$, but were unsuccessful. Maybe there exists some $2$-accessible set $D=\{d_{1},d_{2},d_{3},...\}$ such that $d_{i+1}\geq(2+\delta)d_{i}$? For example, it is not known whether the set of Pell numbers, which satisfy the recurrence $p_{n}=2p_{n-1}+p_{n-2}$, are $2$-accessible or not.
We would also like to determine whether the Fibonacci numbers are $3$-accessible or not. If there existed some $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\{\alpha f_{n}\}\in\left[\varepsilon,\frac{2}{3}\right]$ for all $n\geq 1$, then it would follow from the remark at the conclusion of Theorem 1.4 that $\text{doa}(F)=2$. However, it seems unlikely that such an $\alpha\in\mathbb{R}$ exists, since for every $\alpha\not\in\mathbb{Q}(\sqrt{5})$, there exists infinitely many $n\in\mathbb{N}$ such that $\{\alpha f_{n}\}>2/3$ [3].
It would also be interesting to find a set $D$ which is $2$-accessible (and not $2$-large) that does not contain arbitrarily long monochromatic $D$-diffsequences. We have yet to see a proof of a set being $2$-accessible that does not use Corollary 2.6 or the fact that $D-D:=\{d_{2}-d_{1}:d_{1},d_{2}\in D,d_{2}-d_{1}>0\}$ is accessible for any infinite set $D$.