### v2

# Accessibility of Sparse Sets

Abstract.
A set $D\subseteq\mathbb{N}$ is called $r$-large if every $r$-coloring of $\mathbb{N}$ admits arbitrarily long monochromatic arithmetic progressions $a,a+d,...,a+(k-1)d$ with gap $d\in D$. Closely related to largeness is accessibility; a set $D\subseteq\mathbb{N}$ is called $r$-accessible if every $r$-coloring of $\mathbb{N}$ admits arbitrarily long monochromatic sequences $x_{1},x_{2},...,x_{k}$ with $x_{i+1}-x_{i}\in D$. It is known that if $D\subseteq\mathbb{N}$ is $2$-large, then the gaps between elements in $D$ cannot grow exponentially. In this paper, we show that if $D$ is $2$-accessible, then the gaps between elements in $D$ cannot grow much faster than exponentially.
1. Introduction
An $r$-coloring of a set $A$ is a function $\chi:A\to[r]$. We say a subset $A^{\prime}\subseteq A$ is monochromatic under $\chi$ if $\chi$ is constant over $A^{\prime}$. One of the main goals in the study of Ramsey theory is to find “order” in seemingly “random” structures. For example, van der Waerden’s theorem tells us that given any $r$-coloring of the positive integers, there will exist arbitrarily long monochromatic arithmetic progressions. The theorem places no requirement on the gap (common difference), $d$, of the arithmetic progression – it can be any natural number. With this in mind, in [1], the authors introduce the notion of largeness.
Definition 1.1.
A set $D\subseteq\mathbb{N}$ is called $r$-large if, for every $r$-coloring of $\mathbb{N}$, there exist arbitrarily long monochromatic arithmetic progressions with gap $d\in D$. If $D$ is $r$-large for every $r\in\mathbb{N}$, we say $D$ is large.
In an attempt to learn more about largeness, in [10], the authors study some Ramsey properties of a larger family of sequences including arithmetic progressions.
Definition 1.2.
Let $D\subseteq\mathbb{N}$. A $k$-term sequence of positive integers $x_{1},x_{2},...,x_{k}$ where $x_{i+1}-x_{i}\in D$ is called a $k$-term $D$-diffsequence.
Definition 1.3.
A set $D\subseteq\mathbb{N}$ is called $r$-accessible if, for every $r$-coloring of $\mathbb{N}$, there exist arbitrarily long monochromatic sequences $x_{1},x_{2},...,x_{k}$ with $x_{i+1}-x_{i}\in D$; that is, there exist monochromatic $k$-term $D$-diffsequences for every $k\geq 1$. If $D$ is $r$-accessible for every $r\geq 1$, we say $D$ is accessible.
Equivalently, by a compactness argument, a set $D\subseteq\mathbb{N}$ is $r$-accessible if, for every $k\geq 1$, there exists a least positive integer $\Delta=\Delta(D,k;r)$ such that whenever $[1,\Delta]$ is $r$-colored, there exists a monochromatic $k$-term $D$-diffsequence. The same is true for $r$-largeness. Given a set $D\subseteq\mathbb{N}$, we define the degree of accessibility of $D$, $\text{doa}(D)$, to be the greatest positive integer $r$ such that $D$ is $r$-accessible. If $D$ is $r$-accessible for every $r\in\mathbb{N}$, we write $\text{doa}(D)=\infty.$ Finally, note that it follows immediately from the definitions of largeness and accessibility that any set $D\subseteq\mathbb{N}$ which is $r$-large is automatically $r$-accessible.
It is a consequence of the polynomial van der Waerden theorem [15] that given any polynomial $p(x)\in\mathbb{Q}[x]$ with positive leading term and zero constant term, the set $\{p(n):n\in\mathbb{N}\}\cap\mathbb{N}$ is large (and thus accessible). This gives us a large class of sets which are large. For example, the result tells us that the set $\{n^{k}:n\in\mathbb{N}\}$ is large for every $k\in\mathbb{N}$. Note that the preceding set has natural density 00 and is large, while the set $\{2k+1:k\geq 0\}$ of odd positive integers has natural density $1/2$ and is not even $2$-accessible (color the even positive integers blue and the odd positive integers red). This seems to suggest that the natural density of a set does not give us much information in regards to largeness and accessibility. However, in [1] the authors show that if $D=\{d_{1},d_{2},d_{3},...,\}$ is a set of positive integers satisfying $d_{n+1}\geq 3d_{n}$ for all $n\geq 1$, then $D$ is not $2$-large. In [2], (Question 5.4) the author asks if a similar result holds for $2$-accessibility. Namely: Does there exist an absolute constant $C\geq 2$ such that there is no $2$-accessible set $D=\{d_{1},d_{2},d_{3},...\}$ satisfying $d_{n+1}>Cd_{n}$ for all $n\in\mathbb{N}$? If yes, what is the smallest such $C$? We prove the following theorem:
Theorem 1.4.
Let $r\geq 2$ and let $D=\{d_{1}<d_{2}<d_{3}<\cdots\}$ be a set of positive integers. If there exists a real number $\delta>0$ and $N\in\mathbb{N}$ such that
$d_{n+1}\geq(2+(r-1)^{-1}+\delta)d_{n}$
for all $n\geq N$, then $\text{doa}(D)\leq r-1$.
Letting $r=2$, we immediately obtain the following corollary:
Corollary 1.5.
Let $D=\{d_{1}<d_{2}<d_{3}<\cdots\}$ be a set of positive integers. If there exists a real number $\delta>0$ and $N\in\mathbb{N}$ such that
$d_{n+1}\geq(3+\delta)d_{n}$
for all $n\geq N$, then $D$ is not $2$-accessible.
Remark: We note that since the upper bound $\text{doa}(D)\leq r-1$ of Theorem 1.4 grows with $r$, the result is only interesting for $r=2$ and other small values of $r$. In general, if $d_{n+1}\geq(1+\delta)d_{n}$, then $\text{doa}(D)<\infty$. In particular, a simple argument shows that if $d_{n+1}\geq(2+\delta)d_{n}$ then $\text{doa}(D)\leq 15$. See, for example [11], which studies the chromatic number of graphs $G=(\mathbb{N},E)$, where $xy\in E$ $\iff$ $|x-y|\in D.$ It follows immediately that if $\chi(G)\leq r$ then $\text{doa}(D)\leq r-1$. However, no growth rate condition is sufficient to conclude $\chi(G)\leq 2$ (and thus $\text{doa}(D)=1$). Indeed, take $D=\{(2k)^{n}:n\geq 0\}$. Then, $\{1,2k+1,4k+1,6k+1,...,(2k)^{2}+1\}$ is an odd cycle in $G=(\mathbb{N},E)$, and thus $\chi(G)\geq 3$.
Corollary 1.5 is strikingly similar to the above mentioned growth rate condition for $2$-largeness; the only difference being the added $\delta>0$. The $3$ found in both results seems to be a coincidence, however. As we will see in Section 4, the condition $d_{n+1}\geq 3d_{n}$ in the $2$-largeness result can be lowered to $d_{n+1}\geq(1+\delta)d_{n}$, where $\delta>0$ is any positive constant [6]. On the other hand, the set $\{2^{n}:n\geq 0\}$ is $2$-accessible and has $d_{n+1}\geq 2d_{n}$.
In [8], the authors ask about the degree of accessibility of the set of Fibonacci numbers $F=\{1,2,3,5,8,13,...\}$ and show that $2\leq\text{doa}(F)\leq 5$. Recently, in [16], the author lowers the upper bound, showing $\text{doa}(F)\leq 3$. Their proof uses various properties of the Fibonacci sequence to construct a $4$-coloring of $\mathbb{N}$ that avoids arbitrarily long monochromatic $F$-diffsequences. In this paper, we arrive at the same upper bound of $\text{doa}(F)\leq 3$ using only the growth rate of the Fibonacci sequence.
2. Some Background and Basic Results
First, we note that if $D$ is a finite set, then $D$ is neither $2$-large nor $2$-accessible. To see this, let $n=\max(D)$. Then, the $2$-coloring $\chi:\mathbb{N}\to\{1,2\}$ such that
$\chi(x)=\begin{cases}1&\text{if }1\leq(x\bmod{2n})\leq n\\ 2&\text{otherwise }\end{cases}$
admits no monochromatic $(n+1)$-term $D$-$AP$ and no $(n+1)$-term $D$-diffsequence.
The following result gives us a quite powerful necessary condition for a set $D$ to be $2$-large. A proof can be found in [1].
Theorem 2.1 ([1], Theorem 2.1).
If $D\in\mathbb{N}$ is $2$-large then $D$ contains a multiple of $m$ for every $m\in\mathbb{N}$.
We can also remove a finite number of elements from a $r$-large set and preserve $r$-largeness. This is not in general true for accessibility; for example, the set of odd positive integers $D=\{1,3,5,7,...\}$ is not $2$-accessible but $D\cup\{2\}$ is $3$-accessible [8].
Theorem 2.2 ([1], Lemma 1).
Let $D$ be $r$-large. If $F$ is a finite set, then $D\setminus F$ is $r$-large.
We also have the previously mentioned growth rate condition for $2$-largeness.
Theorem 2.3 ([1], Theorem 2.2).
Let $D=\{d_{1},d_{2},d_{3},...\}$ be a set of positive integers such that $d_{n+1}\geq 3d_{n}$ for all $n\geq 1$. Then $D$ is not $2$-large.
The following tells us that if $D=D_{1}\cup D_{2}$ is large, then at least one of $D_{1}$ or $D_{2}$ is also large. A inductive argument naturally generalizes to $D=\bigcup_{i=1}^{n}D_{i}$ (let $D_{1}=\bigcup_{i=1}^{n-1}D_{i}$ and $D_{2}=D_{n}$).
Theorem 2.4 ([1], Theorem 2.4).
If $D=D_{1}\cup D_{2}$ is large, then either $D_{1}$ or $D_{2}$ is large.
Probably the most general characterization of large sets is a consequence of the polynomial van der Waerden theorem:
Theorem 2.5 ([1], Theorem 3.1).
Let $p$ be a polynomial with integer coefficients, a positive leading term, and $p(0)=0$. Then, $\mathbb{N}\cap\{p(n):n\geq 1\}$ is large.
Remark: In the above theorem, it states that the polynomial need have integer coefficients. This is not necessary; the polynomial need only have rational coefficients. To see why, let $p(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+...+a_{1}x$ be a polynomial with $a_{n},a_{n-1},...,a_{1}\in\mathbb{Q}$. Then, we can write
$a_{n}=\frac{p_{n}}{q_{n}},a_{n-1}=\frac{p_{n-1}}{q_{n-1}},...,a_{1}=\frac{p_{1}}{q_{1}}$
where $\gcd(p_{i},q_{i})=1$. Let $l=\text{lcm}(q_{1},q_{2},...,q_{n}).$ Then, $q(x)=p(lx)$ is a polynomial with integer coefficients and we can apply Theorem 2.4 to $q(x)$.
There are far fewer known results concerning only accessibility. The following is shown in [7]:
Theorem 2.6 ([7], Lemma 2.1).
Let $c\geq 0$ and $r\geq 2$, and let $D\subseteq\mathbb{N}$. If every $(r-1)$-coloring of $D$ yields arbitrarily long monochromatic $(D+c)$-diffsequences, then $D+c$ is $r$-accessible.
Most useful is the case when $r=2$ and $c=0$, in which case we have:
Corollary 2.7.
Let $D\subseteq\mathbb{N}$. If $D$ itself contains arbitrarily long $D$-diffsequences, then $D$ is $2$-accessible.
This corollary gives us the lower bound of $2$ for the degree of accessibility of the Fibonacci numbers $F$. Since $F=\{f_{n}:n\geq 1\}$ and $f_{n}=f_{n-1}+f_{n-2}$, we have that $f_{n}-f_{n-1}=f_{n-2}$ for all $n\geq 3$. Thus, the set of Fibonacci numbers contain infinitely long Fibonacci diffsequences, and are thus $2$-accessible by Corollary 2.7.
Given some subset $D\subseteq\mathbb{N}$, it is often useful to work with a subset of $D$, as opposed to $D$ itself. This next theorem is a generalization of a result that appears in [16] and [9].
Theorem 2.8.
Let $A\subseteq\mathbb{N}$ and let $d\in\mathbb{N}$. Let $B=\left\{a_{i}/d:a_{i}\in A\right\}\cap\mathbb{N}$ and let $C=\{a_{i}\in A:a_{i}\equiv 0\bmod{d}\}$. Then, the following holds:
(i) If $A$ is $rd$-accessible, then $B$ is $r$-accessible.
(ii) If $A$ is $rd$-accessible, then $C$ is $r$-accessible.
Proof.
In both cases, we will prove the contrapositive. First, assume that $B$ is not $r$-accessible. Then, there exists $k\in\mathbb{N}$ and a coloring $\chi:\mathbb{N}\to[r]$ such that $\chi$ admits no monochromatic $k$-term $B$-diffsequence. Now, assume for the sake of contradiction that $A$ is $rp$-accessible. Define the following $rd$-coloring $\chi_{1}:\mathbb{N}\to[d]\times[r]$ such that
$\chi_{1}(x)=\left(x\bmod{d},\chi\left(\frac{x-x\bmod{d}}{d}+1\right)\right).$
Since $A$ is $rd$-accessible, there exists a monochromatic $k$-term $A$-diffsequence $X=\{x_{1},x_{2},...,x_{k}\}$ under $\chi_{1}$. By definition of $\chi_{1}$, every element in $X$ is congruent to $c\bmod{d}$ for some $c\in\{0,1,...,d-1\}$. But then,
$\chi\left(\frac{x_{1}-c}{d}+1\right)=\chi\left(\frac{x_{2}-c}{d}+1\right)=\cdots=\chi\left(\frac{x_{k}-c}{d}+1\right)$
and
$\left(\frac{x_{j+1}-c}{d}+1\right)-\left(\frac{x_{j}-c}{d}+1\right)=\left(\frac{x_{j+1}-x_{j}}{d}\right)\in B$
for all $j\in\{1,2,...,k-1\}$, a contradiction.
Now, assume that $C$ is not $r$-accessible. Then, there exists $k\in\mathbb{N}$ and a coloring $\chi:\mathbb{N}\to[r]$ such that $\chi$ admits no monochromatic $k$-term $C$-diffsequence. Again, assume for the sake of contradiction that $A$ is $rd$-accessible. Define the following $rd$-coloring $\chi_{1}:\mathbb{N}\to[d]\times[r]$ such that
$\chi_{1}(x)=\left(x\bmod{d},\chi(x)\right).$
Since $A$ is $rd$-accessible, there exists a monochromatic $k$-term $A$-diffsequence $X=\{x_{1},x_{2},...,x_{k}\}$ under $\chi_{1}$. By definition of $\chi_{1}$, every element in $X$ is congruent to $c\bmod{d}$ for some $c\in\{0,1,...,d-1\}$. But then,
$\chi(x_{1})=\chi(x_{2})=\cdots=\chi(x_{k})$
and $x_{j+1}-x_{j}\in A$ and $x_{j+1}-x_{j}\equiv 0\bmod{d}$, so $x_{j+1}-x_{j}\in C$ for all $j\in\{1,2,...,k-1\}$, a contradiction.
Theorem 2.8 will be useful for showing $\text{doa}(F)\leq 3$. By letting $F_{E}=\{f_{3n}:n\geq 1\}=\{2,8,34,...\}$, we see that $F_{E}=\{f\in F:f\equiv 0\bmod{2}\}$. Thus, if we can show $\text{doa}(F_{E})=1$, we will have $\text{doa}(F)\leq 3$.
3. Proof of Theorem 1.4 and some Consequences
We begin by giving a proof of our main result, Theorem 1.4. Our proof, and result, is similar in flavor to that of Theorem 2 in [13], and the same sort of nested interval construction can be found in [4]. We define $\{x\}:=x-\lfloor x\rfloor$ to be the fractional part of $x\in\mathbb{R}$. The first part of the proof uses the fact that $d_{n+1}\geq(2+(r-1)^{-1}+\delta)d_{n}$ to construct an $\alpha\in\mathbb{R}$ satisfying $\{\alpha d_{n}\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $n\geq 1$, where $\varepsilon>0$ is some fixed positive constant. The second part shows that this condition on the fractional parts $\{\alpha d_{n}\}$ is sufficient to conclude $D$ is not $r$-accessible.
Proof of Theorem 1.4.
Let $D=\{d_{1}<d_{2}<d_{3}<\cdots\}$ be a set of positive integers such that there exists a real number $\delta>0$ and $N\in\mathbb{N}$ such that $d_{n+1}\geq(2+(r-1)^{-1}+\delta)d_{n}$ for all $n\geq N$. Let $q_{n}=d_{n+N-1}$. Then, $q_{n+1}\geq(2+(r-1)^{-1}+\delta)q_{n}$ for all $n\geq 1$. We construct a sequence of nested intervals $I_{1}\supseteq I_{2}\supseteq I_{3}\supseteq\cdots$ such that
$I_{n}=\left[\frac{z_{n}+\varepsilon_{1}}{q_{n}},\frac{rz_{n}+(r-1)}{rq_{n}}\right],$
where $z_{n}\in\mathbb{N}\cup\{0\}$ and $\varepsilon_{1}>0$ is fixed. Set
$\varepsilon_{1}=\frac{\delta(r-1)}{r\left(2+(r-1)^{-1}+\delta\right)}$
and define $I_{1}=\left[\frac{\varepsilon_{1}}{q_{1}},\frac{r-1}{rq_{1}}\right].$ Assume we have defined $I_{1}\supseteq I_{2}\supseteq\cdots\supseteq I_{k}$ as above. Thus, we have
$I_{k}=\left[\frac{z_{k}+\varepsilon_{1}}{q_{k}},\frac{rz_{k}+(r-1)}{rq_{k}}\right].$
Note that $I_{k}=I^{(1)}_{k}\cup I^{(2)}_{k}$ where
$I^{(1)}_{k}=\left[\frac{z_{k}+\varepsilon_{1}}{q_{k}},\frac{z_{k}+\varepsilon_{1}}{q_{k}}+\frac{1}{q_{k+1}}\right]$
and
$I^{(2)}_{k}=\left[\frac{z_{k}+\varepsilon_{1}}{q_{k}}+\frac{1}{q_{k+1}},\frac{rz_{k}+(r-1)}{rq_{k}}\right].$
Since $|I^{(1)}_{k}|=\frac{1}{q_{k+1}}$, there must exist $z_{k+1}\in\mathbb{N}$ such that $\frac{z_{k+1}}{q_{k+1}}\in I^{(1)}_{k}$. We claim that
$I_{k+1}=\left[\frac{z_{k+1}+\varepsilon_{1}}{q_{k+1}},\frac{rz_{k+1}+(r-1)}{rq_{k+1}}\right]\subseteq I_{k}.$
Since $\frac{z_{k+1}}{q_{k+1}}\in I_{k}^{(1)}$ and $I_{k}=I_{k}^{(1)}\cup I_{k}^{(2)}$, the above inclusion will hold if $|I^{(2)}_{k}|\geq\frac{r-1}{rq_{k+1}}.$ By construction,
$|I^{(2)}_{k}|=\frac{(r-1)-r\varepsilon_{1}}{rq_{k}}-\frac{1}{q_{k+1}}.$
Using the fact that $q_{k+1}\geq(2+(r-1)^{-1}+\delta)q_{k}$, we obtain
$\frac{(r-1)-r\varepsilon_{1}}{rq_{k}}-\frac{1}{q_{k+1}}$ $\geq\frac{\left((r-1)-r\varepsilon_{1}\right)\left(2+(r-1)^{-1}+\delta\right)}{rq_{k+1}}-\frac{1}{q_{k+1}}$
$=\frac{2(r-1)-2r\varepsilon_{1}+1-r\varepsilon_{1}(r-1)^{-1}+\delta(r-1)-r\varepsilon_{1}\delta-r}{rq_{k+1}}$
$=\frac{r-1}{rq_{k+1}}+\frac{\delta(r-1)-2r\varepsilon_{1}-r\varepsilon_{1}(r-1)^{-1}-r\varepsilon_{1}\delta}{rq_{k+1}}$
$=\frac{r-1}{rq_{k+1}}+\frac{\delta(r-1)-\varepsilon_{1}r(2+(r-1)^{-1}+\delta)}{rq_{k+1}}$
$=\frac{r-1}{r_{q+1}},$
where the last equality follows from the definition of $\varepsilon_{1}.$ Now, since the $I_{n}$ are nested and $|I_{n}|\to 0$ as $n\to\infty$, we have $\bigcap_{n=1}^{\infty}I_{n}=\{\alpha\}$. By construction, this $\alpha\in\mathbb{R}$ satisfies $\{\alpha q_{n}\}\in\left[\varepsilon_{1},\frac{r-1}{r}\right]$ for all $n\geq 1$. However, since $I_{1}=\left[\frac{\varepsilon_{1}}{q_{1}},\frac{r-1}{rq_{1}}\right]$, we have $\frac{\varepsilon_{1}}{d_{N}}=\frac{\varepsilon_{1}}{q_{1}}\leq\alpha\leq\frac{r-1}{rq_{1}}=\frac{r-1}{rd_{N}}$. Thus, letting $\varepsilon=\varepsilon_{1}(d_{1}/d_{N})$, we have $\{\alpha d_{n}\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $n\geq 1$.
Consider the $r$-coloring $\chi:\mathbb{N}\to\{1,2,...,r\}$ defined by
$\chi(x)=\begin{cases}1&\text{if }\{\alpha x\}\in\left[0,\frac{1}{r}\right)\\ 2&\text{if }\{\alpha x\}\in\left[\frac{1}{r},\frac{2}{r}\right)\\ \vdots&\\ r&\text{if }\{\alpha x\}\in\left[\frac{r-1}{r},1\right)\end{cases}$
We claim this coloring avoids monochromatic $\left(\lceil 1/r\varepsilon\rceil+1\right)$-term $D$-diffsequences. To see this, assume, for the sake of contradiction, that $X=\{x_{1},x_{2},...,x_{k}\}$ where $k=\left(\lceil 1/r\varepsilon\rceil+1\right)$ is monochromatic, say color $c$. Then, $\{\alpha x_{1}\}\in\left[\frac{c-1}{r},\frac{c}{r}\right)$. We claim
$\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}$
for all $11\leq n\leq k-$ 1. To see why, assume, for the sake of contradiction, that $\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}-1$. By our assumption that $\chi(X)=c$, we have $\frac{c-1}{r}\leq\{\alpha x_{n}\}<\frac{c}{r}$ and $\varepsilon\leq\{\alpha(x_{n+1}-x_{n})\}\leq\frac{r-1}{r}$ for all $11\leq n\leq k-$ 1. Thus, we have
$\{\alpha x_{n+1}\}=\{\alpha(x_{n+1}-x_{n})\}+\{\alpha x_{n}\}-1<\frac{r-1}{r}+\frac{c}{r}-1=\frac{c-1}{r},$
a contradiction with $\chi(x_{m})=c$. Iterating this fact, we obtain
$\{\alpha x_{k}\}=\{\alpha(x_{k}-x_{k-1})\}+\cdots+\{\alpha(x_{2}-x_{1})\}+\{\alpha x_{1}\}\geq(k-1)\varepsilon+\frac{c-1}{r}\geq\frac{c}{r}.$
Thus, $\chi(x_{k})\not=c$, a contradiction. It follows that $\text{doa}(D)\leq r-1.$
Remark: From the proof of Theorem 1.4, we obtain a new way of showing a set $D$ is not $r$-accessible. Namely: if there exists some $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\{\alpha d\}\in\left[\varepsilon,\frac{r-1}{r}\right]$ for all $d\in D$, then $D$ is not $r$-accessible. The coloring in the proof of Theorem 1.4 can equivalently be defined as
$\chi(x)=r+\sum_{i=1}^{r-1}\left\lfloor\{\alpha x\}-\frac{i}{r}\right\rfloor=1+\sum_{i=1}^{r-1}\left\lfloor\alpha x+\frac{i}{r}\right\rfloor-\lfloor\alpha x\rfloor.$
Using the fact that the set of even Fibonacci numbers satisfy the recurrence $e_{n}=4e_{n-1}+e_{n-2}$ for all $n\geq 3$ (with $e_{1}=2$ and $e_{2}=8$), it follows by Corollary 1.5 that $\text{doa}(F_{E})=1$. Thus, by the comment following the proof of Theorem 2.8, we have $\text{doa}(F)\leq 3$. However, as seen in the following corollary, we can actually calculate an explicit upper bound for the longest monochromatic $F$-diffsequence:
Corollary 3.1.
Let $F_{E}=\{f_{3n}:n\geq 1\}=\{2,8,34,...\}$ be the set of even Fibonacci numbers. Then $\text{doa}(F_{E})=1$.
Proof.
Let $\phi=(1+\sqrt{5})/2$, $\varphi=(1-\sqrt{5})/2$, and $\alpha=1/\sqrt{5}.$ Then,
$f_{n}=\alpha\phi^{n}-\alpha\varphi^{n}$
for all $n\geq 0$. We note that
$\phi f_{n}=f_{n+1}-\varphi^{n}.$
Thus,
$(1+\phi)f_{n}=f_{n}+f_{n+1}-\varphi^{n}.$
We will use the fact that $f_{3n}+f_{3n+1}\equiv 1\bmod{4}.$ This is true for $n\in\{0,1\}$. Now, assume that $k\geq 2$ and $f_{3l}+f_{3l+1}\equiv 1\bmod{4}$ for all $l<k$. Then, $f_{3k}+f_{3k+1}\equiv f_{3(k-2)}+f_{3(k-2)+1}\equiv 1\bmod{4}$ since $f_{n}\equiv f_{n+6}\bmod{4}$ (the Fibonacci sequence is periodic modulo $m$ for every $m$; when $m=4$, the period is $6$). Hence, it follows that
$\left\{\frac{f_{3n}+f_{3n+1}}{4}\right\}=\frac{1}{4}$
for $n\geq 1$. For $n\geq 1$, we also have
$-0.04<-\frac{1}{4}\left(\frac{1-\sqrt{5}}{2}\right)^{4}\leq-\frac{\varphi^{3n}}{4}\leq-\frac{1}{4}\left(\frac{1-\sqrt{5}}{2}\right)^{3}<0.06$
Thus,
$0.310.21<\left\{(1+\phi)\frac{f_{3n}}{4}\right\}<$ 0.31
for all $n\geq 1$ and it follows by the remark following the proof of Theorem 1.4 that $\text{doa}(F_{E})=1$. In fact, the longest monochromatic $F_{E}$-diffsequence under $\chi$ is bounded above by $\lceil 1/2(0.21)\rceil+1=4$.
We conclude this section by noting there seems to be a relationship between the density in $[0,1]$ of sequences ($(\{\alpha d_{n}\})_{n\geq 1}$ where $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ and the degree of accessibility of $D$. We wonder if the following holds (see also [6], Conjecture 1):
Question 1. Let $D=\{d_{1},d_{2},d_{3},...\}\subseteq\mathbb{N}$ be such that every (eventually) periodic $2$-coloring of $\mathbb{N}$ admits arbitrarily long monochromatic $D$-diffsequences. If ($(\{\alpha d_{n}\})_{n\geq 1}$ is dense in $[0,1]$ for every $\alpha\in\mathbb{R}\setminus\mathbb{Q}$, is $D$ $2$-accessible?
If the above is true, then letting $p_{n}$ denote the $nth$ prime number and using the fact that ($(\{\alpha p_{n}\})_{n\geq 1}$ is dense in $[0,1]$ for any $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ [14], it would follow from the remark at the end of [10] that the primes are $2$-accessible.
4. A Small Digression Into Largeness
It was mentioned in the introduction that the $d_{n+1}\geq 3d_{n}$ condition in Theorem 2.3 could be weakened to $d_{n+1}\geq(1+\delta)d_{n}$, where $\delta$ is any positive constant [6]. Since the argument is very similar to that used in Theorem 1.4, we include a proof of the result in this section. We also give a $2$-coloring of $\mathbb{N}$ that does not admit monochromatic $11$-term arithmetic progressions with Fibonacci differences.
We define $||x||:=\min(\{x\},1-\{x\})$, that is, the distance from $x$ to the nearest integer. We require the following result on sequences with exponential growth (see [11] for subsequent work).
Theorem 4.1 ([12]).
Let $(d_{n})_{n\geq 1}$ be a sequence of positive integers such that
$d_{n+1}\geq(1+\delta)d_{n}$
for all $n\geq 1$. Then, there exists $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\|\alpha d_{n}\|\geq\varepsilon$ for all $n\geq 1$.
The main component of the proof of Theorem 4.3 relies on the following result. A proof appears in ([6], Theorem 2).Our argument is practically the same; however, we use slightly different notation in our proof.
Theorem 4.2.
Let $D\subseteq\mathbb{N}$. If there exists an $\alpha\in\mathbb{R}$ and $\varepsilon>0$ such that $\|\alpha d\|\geq\varepsilon$ for all $d\in D$, then $D$ is not $2$-large.
Proof.
Since $||\alpha d||\geq\varepsilon$ for all $d\in D$, we have $\{\alpha d\}\in[\varepsilon,1-\varepsilon]$ for all $d\in D$. Now, we define the $2$-coloring $\chi:\mathbb{N}\to\{1,2\}$ by
$\chi(x)=\begin{cases}1:\{\alpha x\}\in\left[0,\frac{1}{2}\right)\\ 2:\{\alpha x\}\in\left[\frac{1}{2},1\right).\end{cases}$
Since $\{\alpha d\}\in[\varepsilon,1-\varepsilon]$ for all $d$, for any monochromatic arithmetic progression with $d\in D$ under $\chi$, we would have either $\{\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$ or $\{\alpha d\}\in\left[\frac{1}{2},1-\varepsilon\right]$. Now, if $\{\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$, then the same argument used in the latter part of Theorem 1.4 shows us that we cannot have a monochromatic $(\lfloor 1/2\varepsilon\rfloor+1)$-term arithmetic progression with gap $d$. If $\{\alpha d\}\in\left[\frac{1}{2},1-\varepsilon\right]$, then $\{-\alpha d\}\in\left[\varepsilon,\frac{1}{2}\right]$. Thus, the same argument used backwards shows us that we cannot have a monochromatic $(\lfloor 1/2\varepsilon\rfloor+1)$-term arithmetic progression with gap $d$. Thus, $D$ is not $2$-large. ∎
Remark: Theorem 4.2 implies Theorem 2.1 since if $D\subseteq\mathbb{N}$ contains no multiple of $m\in\mathbb{N}$, then $||\frac{1}{m}d||\geq\frac{1}{m}$ for all $d\in D$.
With Theorems 4.1 and 4.2, we immediatly obtain the following:
Theorem 4.3.
Let $D=\{d_{1},d_{2},d_{3},...\}$ be a set of positive integers. If there exists a $\delta>0$ such that
$d_{n+1}\geq(1+\delta)d_{n}$
for all $n\geq 1,$ then $D$ is not $2$-large.
It follows immediately from Theorem 4.3 that the set of Fibonacci numbers $F$ are not $2$-large. However, using only Theorem 4.2 and some properties of the Fibonacci sequence we can obtain an upper bound on the longest monochromatic arithmetic progression with gap $f\in F$:
Corollary 4.4.
The set $F=\{1,2,3,5,8,...\}$ of Fibonacci numbers is not $2$-large.
Proof.
Let $\phi=(1+\sqrt{5})/2$, $\varphi=(1-\sqrt{5})/2$, and $\alpha=1/\sqrt{5}.$ Then,
$f_{n}=\alpha\phi^{n}-\alpha\varphi^{n}$
for all $n\geq 0$. We note that
$\phi^{-1}f_{n}=f_{n-1}-\varphi^{n}$
and
$\phi f_{n}=f_{n+1}-\varphi^{n}.$
Thus,
$\sqrt{5}f_{n}=(\phi^{-1}+\phi)f_{n}=f_{n-1}+f_{n+1}-2\varphi^{n}.$
We will use the fact that $f_{n-1}+f_{n+1}\not\equiv 0\bmod{8}$ for all $n\geq 1$. This can be easily verified by checking the first few cases and then noting that $f_{n}\equiv f_{n+12}\bmod{8}.$ Direct computation gives us
$\left\{\frac{\sqrt{5}f_{1}}{8}\right\}=\left\{\frac{\sqrt{5}f_{2}}{8}\right\}\approx 0.28$
Then, for $n\geq 3$, we note that
$-0.04<-\frac{1}{4}\left(\frac{1-\sqrt{5}}{2}\right)^{4}\leq-\frac{\varphi^{n}}{4}\leq-\frac{1}{4}\left(\frac{1-\sqrt{5}}{2}\right)^{3}<0.06.$
Since
$\left\{\frac{f_{n-1}+f_{n+1}}{8}\right\}\in\left\{\frac{1}{8},\frac{2}{8},...,\frac{7}{8}\right\},$
we have
$0.9350.085<\left\{\frac{\sqrt{5}f_{n}}{8}\right\}<$ 0.935 .
Thus, $\left\|\frac{\sqrt{5}}{8}f\right\|\geq 0.06$ for all $f\in F$ and $F$ is not $2$-large by Theorem 4.2. In fact, the longest monochromatic arithmetic progression with gap $f\in F$ under $\chi$ is bounded above by $\lceil 1/2(0.06)\rceil+1=10$.
5. Some Concluding Remarks
The condition $d_{n+1}\geq(1+\delta)d_{n}$ in Theorem 4.3 is tight. On the other hand, we are not sure whether the condition $d_{n+1}\geq(3+\delta)d_{n}$ in Corollary 1.5 is the best possible. We hoped to weaken the condition to $d_{n+1}\geq(2+\delta)d_{n}$, but were unsuccessful. When $d_{n+1}\geq\lambda d_{n}$ with $\lambda\in(2,3]$, the nested interval construction in Theorem 1.4 breaks down. A somewhat related problem ([3], Problem $5$) asks weather or not there exists an $\alpha>2$ such that $\{\zeta\alpha^{n}\}\geq 1/2$ infinity often for all $\zeta\not=0$.
Maybe there exists some $2$-accessible set $D=\{d_{1},d_{2},d_{3},...\}$ such that $d_{n+1}\geq(2+\delta)d_{n}$? For example, it is not known whether the set of Pell numbers, which satisfy the recurrence $p_{n}=2p_{n-1}+p_{n-2}$, are $2$-accessible or not.
We would also like to determine whether the Fibonacci numbers are $3$-accessible or not. We cannot apply Theorem 1.4 to conclude they are not $3$-accessible since for any $\varepsilon>0$ and $\alpha\in\mathbb{R}$ it is impossible to have $\{\alpha f_{n}\}\in\left[\varepsilon,\frac{2}{3}\right]$ for all $n$. See [5], which proves that for every $\alpha\not\in\mathbb{Q}(\sqrt{5})$, there exists infinitely many $n\in\mathbb{N}$ such that $\{\alpha f_{n}\}>2/3$. If $\alpha\in\mathbb{Q}(\sqrt{5})$ then (using the fact that $\sqrt{5}f_{n}=f_{n-1}+f_{n+1}-2\varphi^{n}$) a fairly straightforward argument can show that for any $\varepsilon>0$, it is impossible to have $\{\alpha f_{n}\}\in\left[\varepsilon,\frac{2}{3}\right]$ for all $n$.
It would also be interesting to find a set $D$ which is $2$-accessible (and not $2$-large) that does not contain arbitrarily long monochromatic $D$-diffsequences. We have yet to see a proof of a set being $2$-accessible that does not use Corollary 2.7 or the fact that $D-D:=\{d_{2}-d_{1}:d_{1},d_{2}\in D,d_{2}-d_{1}>0\}$ is accessible for any infinite set $D$.